# Diffusion Forcing for Multi-Agent Interaction Sequence Modeling

### v1

@@ 4.3 Qualitative Evaluation
For qualitative results we refer the reader to our supplementary video where we show various downstream applications of our model. Our qualitative results demonstrate that our model generates diverse, long, and natural motion sequences, capturing dyadic and polyadic movement spanning large spaces. In Figures 1 and 4, we show inpainting task and in-betweening, and ultra-long generation over multiple samples illustrating the diversity, as well as 3 and 4 agent generations.
@@ 5 Conclusion
We present MAGNet,Multi-Agent Diffusion Forcing Transformer , a unified multi-agent autoregressive generative model. Our approach integrates diverse motion generation capabilities—dyadic motion prediction, partner inpainting and prediction, ultra-long motion synthesis, motion control, and scalable multi-agent generation—within a single model. By representing per-person poses through a VQ-VAE codebook and modeling intra- and inter-agent dynamics with the Diffusion Forcing Transformer, our method enables coherent, long-horizon, and diverse multi-agent motion synthesis. Please see the appendix on discussion of limitations. We hope our flexible architecture can serve as a foundation for polyadic agent generation. Future work includes swarm-like generation and scaling to hundreds of agents that interact in a socially plausible manner and exploring other conditioning signals such as text.
@@ 6 Acknowledgments
Our supplementary material also includes a video file that show motions generated by our model for different datasets (dancing, interaction, sports), ultra-long videos, and generation of multiple people.
@@ A.1 Data
  • Contact Sports: We use the DuoBoX[cen2025readytoreact] dataset to represent high-contact motion sequences.
  • Dance: We include several motion capture dance datasets, namely ReMoCap (LindyHop)[ghosh2024remos] and DD100 [siyao2024duolando].
  • Day-to-Day Interactions: For everyday activities, we consider Embody 3D [mclean2025embody], AMASS [amass2019], and Inter-X [xu2024inter].
Among these, AMASS primarily contains single person (unary) motion, Embody 3D features one to four people interactions and the remaining datasets consist of two person (dyadic) interactions. Table A.1 summarizes the key statistics of these datasets, including frame rates, number of clips, subjects, actions, total frames, and approximate total duration.
@@
DD100 [siyao2024duolando] 30 167 5 10 genres 350.4K 3.24 hrs
DuoBox [cen2025readytoreact] 120 116 3 1 913.8K 2.1 hrs
ReMoCap (LindyHop) [ghosh2024remos] 50 8 4 1 174.2K 56 min
Inter-X [xu2024inter] 120 11,388 89 40 8,071.8K 18.68 hrs
AMASS [amass2019] varies 11,265 344 - - 40 hrs
Embody 3D (1 person) [mclean2025embody] 30 16,965 77 597 7,791.6K 72 hrs
Embody 3D (2 people) [mclean2025embody] 30 66 4 24 577.7K 5.3 hrs
Embody 3D (3 people) [mclean2025embody] 30 35 3 18 273.0K 2.5 hrs
Embody 3D (4 people) [mclean2025embody] 30 612 70 56 5,169.7K 47 hrs
@@ A.3 Model Architecture Details
Module Component Input shape Operation / structure Main hyperparameters
VQ-VAE Encoder $(T,P,D_{\text{in}}+D_{\text{cond}})$ 1D Conv + 1D ResNet stack 2 layers, hidden dim $d_{\text{vq}}=512$
Codebook $(T^{\prime},P,d_{\text{vq}})$ Vector quantization Codebook size $K=1024$, embedding dim $d_{\text{vq}}$
@@
Person emb. $(T^{\prime},P,d_{\text{emb}})$ Learned embedding $d_{\text{emb}}=256$
Input emb. $(T^{\prime},P,D_{\text{m}}+2d_{\text{emb}})$ Linear + LayerNorm 3 layers, hidden dim $d=512$, $D_{\text{m}}+2d_{\text{emb}}\rightarrow d$
Time emb. $(T^{\prime},P,d)$ RoPE Project back to hidden dim $d$
@@
Output head $(T^{\prime},P,d)$ Linear projection 1 layer, $d\rightarrow D_{\text{m}}$
Table A.2: Network configuration. Network configuration used in all experiments, where $T$ denotes the number of timesteps, $T^{\prime}=T/4$ the number of DFoT nodes, and $P$ the number of agents.
@@ A.4.2 Inference Speed
Table A.3 demonstrates that MAGNet runs at up to 56 FPS, this means MAGNet can generate one future motion frame in under 18 ms for both partner and dyadic future prediction tasks, which significantly faster than competing methods. This inference speed experiment was conducted on a single NVIDIA RTX A6000 gpu. In the Partner Inpainting task, we achieve 54 FPS (vs. 49 FPS for Duolando and just 1 FPS for ReMoS). For Partner Prediction, our method provides a $\sim 3.5\times$ speedup over Ready-to-React (56 vs. 16 FPS), and for Dyadic future prediction we are $\sim 6.8\times$ faster (54 vs. 8 FPS). These results highlight MAGNet’s ability to deliver real-time performance while maintaining SoTA level motion quality and robust motion generation capabilities.
@@
A.4.2.1 Motion History Guidance
@@
A.4.3 Other Sampling Strategies
@@
Motion Control. In the motion control task, Agent A’s next action ($B_{t}$) is predicted based on all of its past motion and the partner’s current motion ($A_{t}$), enabling direct and adaptive control of Agent B’s behavior in response to the partner’s movements:
$P(B_{t}\mid A_{0:t},B_{0:t-1})$ (26)
This formulation allows Agent A to serve as a motion controller of Agent B’s motion, facilitating responsive and coordinated interaction (see Fig A.2).

### v2

@@ 4.3 Qualitative Evaluation
Our qualitative results corroborate the quantitative findings. In Figure 5(a), our partner predictions maintain plausible boxing exchanges over 150 frames and generate timely reactive motion such as counter-punches in response to the context agent’s strikes, while R2R produces less reactive motions. Figure 5(b) highlights the long-horizon advantage: MAGNet sustains coherent two-person boxing well beyond the 139-frame ground truth (up to 1800 frames), whereas R2R’s agents gradually drift apart and the interaction breaks down. Figure 6 demonstrates polyadic generation on Embody3D with three and four agents using the same trained model, producing coordinated group interactions without architectural changes. Figure 4 illustrates the breadth of tasks handled by a single model: boxing, casual interaction, polyadic generation (P=4), and in-betweening. We encourage viewing the supplementary video, where temporal dynamics and inter-agent coordination are best appreciated.
@@ 5 Conclusion
We present MAGNet, a unified framework for multi-agent motion generation that subsumes dyadic and polyadic future prediction, partner inpainting, ultra-long synthesis, and agentic generation—tasks that previously required separate, specialized models. This is enabled by two core design choices: a pairwise relative transform representation that generalizes to arbitrary numbers of agents by construction, and diffusion forcing over multi-agent tokens that implicitly learns all conditional distributions over any subset of agents and frames. Without text conditioning, MAGNet is on-par and outperforms specialized baselines across multiple motion generation tasks: we improve FD by 89% on long-horizon generation over R2R [2], achieve the lowest FD on DD100 partner inpainting (0.05 vs. 18.18 for Duolando), and scale gracefully from two to four agents on Embody3D where retrieval baselines degrade. We discuss limitations in the appendix. We hope our unified formulation can serve as a foundation for broader multi-agent generation, with future work targeting swarm-scale scenarios involving hundreds of socially interacting agents.
@@ 6 Acknowledgments
Our supplementary material also includes a video file demonstrating motions generated by our model across different datasets (dancing, interaction, sports), as well as ultra-long sequences and multi-person generation.
@@ A.1 Data
  • Contact Sports: We use the DuoBoX[2] dataset to represent high-contact motion sequences.
  • Dance: We include several motion capture dance datasets, namely ReMoCap (LindyHop)[9] and DD100 [26].
  • Day-to-Day Interactions: For everyday activities, we consider Embody 3D [20], and Inter-X [31].
Table A.1 summarizes the key statistics of these datasets, including frame rates, number of clips, subjects, actions, total frames, and approximate total duration.
@@
DD100 [26] 30 167 5 10 genres 350.4K 3.2 hrs
DuoBox [2] 120 116 3 1 913.8K 2.1 hrs
ReMoCap (LindyHop) [9] 50 8 4 1 174.2K 0.9 hrs
Inter-X [31] 120 11,388 89 40 8,071.8K 18.7 hrs
Embody 3D (2 people) [20] 30 66 4 24 577.7K 5.3 hrs
Embody 3D (3 people) [20] 30 35 3 18 273.0K 2.5 hrs
Embody 3D (4 people) [20] 30 612 70 56 5,169.7K 47.9 hrs
@@ A.3 Model Architecture Details
Module Component Input shape Operation Hyperparameters
VQ-VAE Encoder $(T,P,D_{\text{in}}+D_{\text{cond}})$ 1D Conv + ResNet stack 2 layers, hidden dim $d_{\text{vq}}=512$
Codebook $(T^{\prime},P,d_{\text{vq}})$ Vector quantization $K=1024$ codes, dim $d_{\text{vq}}$
@@
Agent emb. $(T^{\prime},P,d_{\text{emb}})$ Learned embedding $d_{\text{emb}}=256$
Input proj. $(T^{\prime},P,D_{\text{m}}+2d_{\text{emb}})$ Linear + LayerNorm 3 layers, $D_{\text{m}}+2d_{\text{emb}}\rightarrow d=512$
Time emb. $(T^{\prime},P,d)$ RoPE Applied to hidden dim $d$
@@
Output head $(T^{\prime},P,d)$ Linear projection $d\rightarrow D_{\text{m}}$
Table A.2: Network configuration. $T$ is the input sequence length where $T^{\prime}=T/\omega$ is the number of DFoT tokens ($\omega=4$) after temporal compression, $P$ the number of agents, $D_{\text{in}}$ the per-agent pose dimension, $D_{\text{cond}}$ the conditioning dimension, and $D_{\text{m}}=d_{\text{vq}}+4\times 9\times P$ the full motion token dimension including pairwise transforms.
@@ A.4.2 Inference Speed
Table A.3 demonstrates that MAGNet runs at up to 56 FPS, this means MAGNet can generate one future motion frame in under 18 ms for both partner and dyadic future prediction tasks, which significantly faster than competing methods. This inference speed experiment was conducted on a single NVIDIA RTX A6000 gpu with no smoothing guidance. In the Partner Inpainting task, we achieve 54 FPS (vs. 49 FPS for Duolando and just 1 FPS for ReMoS). For Partner Prediction, our method provides a $\sim 3.5\times$ speedup over Ready-to-React (56 vs. 16 FPS), and for Dyadic future prediction we are $\sim 6.8\times$ faster (54 vs. 8 FPS). These results highlight MAGNet’s ability to deliver real-time performance while maintaining SoTA level motion quality and robust motion generation capabilities.
@@
A.5 Joint Generation of Social Interaction
We evaluate MAGNet on the task of joint future prediction across two social interaction datasets: InterX [31], a dyadic interaction dataset, and Embody3D [20], which contains scenes with 2–4 interacting people. For Embody3D, we train a single unified model on the full multi-person training set and evaluate separately on the 2-, 3-, and 4-person subsets, demonstrating our framework’s ability to handle a variable number of agents without requiring separate models. Results are reported in Tab. A.4. On InterX, MAGNet achieves strong generation quality with an FID of 0.210 and diverse outputs (DIV of 2.911), while maintaining low reconstruction errors (MPJPE of 0.475, MPJVE of 0.013). On Embody3D, the model generalizes well across varying group sizes. Notably, the 3-person subset yields the lowest FID (0.293) and MPJPE (0.526), suggesting that the model captures group dynamics most effectively at this scale, possibly due to a favorable balance of interaction complexity and data availability. The 2- and 4-person subsets exhibit higher FID scores (0.298 and 0.293, respectively), which may reflect greater distributional variability in dyadic scenes and increased combinatorial complexity in 4-person interactions. Across all subsets, the model maintains consistently low MPJVE values ($\leq$0.009), indicating temporally smooth and physically plausible generated motions. These results highlight the flexibility of our unified multi-agent formulation, which accommodates varying numbers of interacting agents within a single model while producing high-quality, diverse, and physically coherent social interactions.
Dataset Joint Future Prediction
FID DIV MI FS IP MPJPE MPJVE
InterX 2 people 0.210 2.911 0.130 0.074 0.093 0.475 0.013
Embody3D 2 people 0.298 6.219 0.093 0.207 0.023 0.561 0.011
3 people 0.293 6.313 0.004 0.257 0.043 0.526 0.010
4 people 0.590 3.018 0.058 0.237 0.050 0.546 0.009
Table A.4: Evaluation on Social Interaction Dataset. Our model is evaluated on joint future prediction on InterX and Embody3D. For Embody3D, we train a single model on the full multi-person set (2–4 people) and report metrics separately on the 2-, 3-, and 4-person subsets.
A.6 Additional Ablations
A.6.1 Design Choice Ablations
Table A.5 ablates our positional encoding design. Agent PE informs the model which agent a token belongs to, while Time PE encodes when it occurs in the sequence. Removing Agent PE degrades FD by $11\times$ (0.560 vs. 0.052) and foot skating by $2.5\times$, indicating the model can no longer distinguish between agents and produces spatially incoherent motion. Removing Time PE has an even larger impact on motion quality (FD: 2.032), as the model loses temporal ordering and generates implausible transitions. Removing both yields the worst velocity estimation (MPJVE: 0.094) despite artificially high diversity, suggesting the model collapses to producing varied but unstructured motion. Our full model achieves the best scores across all metrics except diversity, which we attribute to the stronger coordination constraints learned when both embeddings are present.
[…]
w/o Agent PE 0.560 14.029 0.148 1.078 0.382 0.810 0.075
w/o Time PE 2.032 15.062 0.365 0.522 0.363 0.828 0.061
w/o Agent & Time PE 0.920 15.081 0.032 0.974 0.370 0.835 0.094
Full model (Ours) 0.052 9.572 0.124 0.423 0.116 0.641 0.034
Table A.5: Ablation of positional encodings. Removing Agent or Time PE degrades motion quality (FD) by $11\times$ and $39\times$ respectively. Our full model achieves the best quality and coordination across all metrics. The higher DIV in ablated variants reflects unstructured motion rather than meaningful diversity.
A.6.2 Noise Sampling Strategy Ablations
Table A.6 compares two noise sampling strategies for the joint future prediction task on DuoBox. In Joint Sampling, both agents at each timestep share the same noise level, preserving spatiotemporal correlations during denoising. In Independent Sampling, each agent’s token receives an independently sampled noise level. Joint Sampling substantially outperforms Independent Sampling on FD ($0.052$ vs. $0.716$), diversity ($9.572$ vs. $1.804$), and interaction quality (IP: $0.116$ vs. $0.231$). Independent Sampling achieves lower foot skating and MPJPE, suggesting it produces kinematically conservative motion—but at the cost of diversity and expressivity. We attribute this to the decorrelated noise breaking the joint structure between agents: each agent denoises plausible but overly safe motion independently, resulting in muted interactions. We use Joint Sampling for all joint prediction experiments.
[…]
Independent (per-token) 0.716 1.804 0.155 0.223 0.231 0.608 0.021
Joint (Ours) 0.052 9.572 0.124 0.423 0.116 0.641 0.006
Table A.6: Joint vs. independent noise sampling on DuoBox. Joint sampling preserves inter-agent correlations during denoising, yielding substantially better distribution quality (FD) and diversity while maintaining strong coordination. Independent sampling produces kinematically conservative but less expressive motion.
A.7 Motion History Guidance
@@
A.8 Smoothing Guidance
During inference the denoiser $f_{\boldsymbol{\phi}}$ iteratively refines the noisy token sequence $\mathbf{M}^{(k)}$, producing at each step a clean-motion estimate $\hat{\mathbf{M}}_{\mathbf{0}}=f_{\boldsymbol{\phi}}(\mathbf{M}^{(k)},\boldsymbol{k}_{\text{seq}})$. Each denoising step sees only the current noisy state and predicts the clean tokens independently—nothing forces adjacent token timesteps to agree on velocity or acceleration. The result can be a motion that is plausible per-frame but jittery when played back.
Smoothness objective.
We inject a preference for temporal coherence directly into the sampling process. At each denoising step we inspect the predicted clean sequence $\hat{\mathbf{M}}_{\mathbf{0}}$ for roughness via the discrete second-order finite difference (acceleration) along the token-time axis. For the $p$-th agent at token timestep $i$:
$\Delta^{2}\hat{\mathbf{m}}^{p}_{0}[i]\;=\;\hat{\mathbf{m}}^{p}_{0}[i-1]\;-\;2\,\hat{\mathbf{m}}^{p}_{0}[i]\;+\;\hat{\mathbf{m}}^{p}_{0}[i+1],\\ i\in\{2,\dots,T^{\prime}-1\},$ (12)
where $T^{\prime}=T/\omega$ is the number of token timesteps after VQ-VAE temporal compression. This quantity is zero for constant-velocity motion and large wherever the trajectory has kinks. We aggregate it into a scalar roughness objective over all $P$ agents:
$\mathcal{L}_{\mathrm{smooth}}\;=\;\sum_{p=1}^{P}\sum_{i=2}^{T^{\prime}-1}\left\|\,\Delta^{2}\hat{\mathbf{m}}^{p}_{0}[i]\,\right\|^{2}.$ (13)
Guided sampling step.
The gradient $\nabla_{\mathbf{M}^{(k)}}\mathcal{L}_{\mathrm{smooth}}$ indicates the direction in noise-space that would most reduce the predicted roughness. We obtain it by backpropagating Eq. (13) through the denoiser $f_{\boldsymbol{\phi}}$ and modify the standard denoising update:
$\mathbf{M}^{(k)}\;\leftarrow\;\mathbf{M}^{(k)}\;-\;\lambda\,\nabla_{\mathbf{M}^{(k)}}\mathcal{L}_{\mathrm{smooth}},$ (14)
where $\lambda$ controls the guidance strength. This steers the sampling trajectory toward solutions that are not only plausible under the learned distribution but also temporally coherent, without any retraining.
Relation to classifier guidance.
This follows the same principle as classifier guidance in image diffusion [6]: an external signal—there a classifier, here a hand-crafted smoothness objective—steers generation at inference time via $\nabla_{\mathbf{M}^{(k)}}$ without modifying the pretrained denoiser $f_{\boldsymbol{\phi}}$. A single model produces either raw samples or temporally smooth ones, controlled entirely by $\lambda$.
Figure A.3: Sampling strategies supported by a single MAGNet model at inference time. The top row shows joint prediction, full sequence denoising, partner inpainting, and in-betweening. The bottom row shows partner prediction, motion control, agentic generation (synchronous), and agentic generation (asynchronous). Red and blue represent Agent A and B, with saturation indicating noise level across denoising steps. Dashed outlines mark empty (unavailable) token positions.
A.9 Other Sampling Strategies
Additional sampling strategies not discussed in the main paper include In-Betweening and Motion Control, which we describe below. Figure A.3 illustrates all the sampling strategies we have explored.
@@ A.9 Other Sampling Strategies
Motion Control. In the motion control task, Agent A’s next action ($A_{t}$) is predicted based on all of its past motion and the partner’s current motion ($B_{t}$), enabling direct and adaptive control of Agent A’s behavior in response to the partner’s movements:
$P(A_{t}\mid A_{0:t-1},B_{0:t})$ (17)
This formulation allows Agent B to serve as a motion controller of Agent A’s motion, facilitating responsive and coordinated interaction (see Fig A.2).